Please help me solve this equation: { { { -2x^2 - 4x - 2 } } }.
Answer by arunpaul(104)

-2x^2 - 4x - 2 =0
or x^2+2x+1 =0 (deviding each side by -2)
or (x+1)^2 =0
or x+1 =0
or x =-1

or

-2x^2 -4x -2 =0
or -2x^2 -2x -2x -2 =0
or -2x(x+1)-2(x+1)=0
or (x+1)(-2x-2) =0
in either case x =-1 (ans)